OpenStudy (anonymous):

Simplify the radical expression √72x^2

OpenStudy (anonymous):

A. 36√x^2 B. 6x√2 C. x√36 D. 6√2x^2

TheSmartOne (thesmartone):

\(\bf\huge~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 @JudgementDay

OpenStudy (anonymous):

O_O that's some big letters.

TheSmartOne (thesmartone):

Haha, so lets solve your question shall we? xD

TheSmartOne (thesmartone):

Do you know this property? \(\sf\Large\sqrt{ab}=\sqrt{a} \times \sqrt{b}\) We can use that to make your question \(\sf\Large\sqrt{72x^2}=\sqrt{72} \times \sqrt{x^2}\)

TheSmartOne (thesmartone):

Well since you are gone, here is a hint to help you solve your question: Hint: \(\sf\Large \sqrt{72} = \sqrt{36\times2} = \sqrt{36}\times\sqrt{2}\)